OpenStudy (anonymous):

Help please, i will give you a medal. 4. A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

I have a "shortcut". Sorry it is a cheat, use the website "wolframalpha.com"

OpenStudy (anonymous):

Once you graph it will probably show you a chart with values of the graph like min, max, vertex,etc

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Your welcome :)

OpenStudy (anonymous):

do you mind helping with one more?

OpenStudy (jdoe0001):

\(\large |x - 0.525| = 0.01\implies \begin{cases} +(x - 0.525) = 0.01\\ -(x - 0.525) = 0.01 \end{cases}\) that should give a min and max values for "x"